Suppose that E and F are two events and that P(E) = 8 and P(FE)- 4. What is P(E and F)?
Answer #1

Ans) Given data PLE) = 0.8, P(Fle) -0.4 P(EMF) = P(E)xplile) 70-8X0.4 =0.32